OpenStudy (anonymous):

e^ infinity = 0 is it true If yes how, prove it

OpenStudy (anonymous):

That's undefined

OpenStudy (anonymous):

no and never

OpenStudy (anonymous):

Yes, but you can take the limit. What is \(\lim_{x\rightarrow \infty} e^x\)?

OpenStudy (anonymous):

And of course it diverges to infinity as x diverges to infinity.

OpenStudy (anonymous):

Maybe you are thinking of negative infinity?

OpenStudy (anonymous):

No. It's just another way to prove \(e^\infty \not = 0\).

OpenStudy (anonymous):

\[e ^∞=∞. For all x>1 we have \lim n→∞ x ^n=∞. Think of \it \in terms of a simple example 2 ^1, 2 ^2, 2 ^3, and so on. It \gets bigger and bigger.\]

OpenStudy (anonymous):

Look it is not defined to begin with. So all of that math is just wrong to begin with. You can express it as a limit but infinity itself is not part of the real number field.

OpenStudy (tkhunny):

\(e^{\infty}\) is not a meaningful expression. "\(= \infty\)" normally should be read "increases without bound". It's not really equality in even an informal sense.